OpenStudy (anonymous):

8/9-a^2 divided by 4a^2-4a-24/a^2-6a+9=

OpenStudy (anonymous):

@Hero i need help

OpenStudy (anonymous):

@KatieMO

OpenStudy (anonymous):

@amistre64 (Moderator)

OpenStudy (anonymous):

Simplify both parts first\[\frac{ 4a^2-4a-24 }{ a^2-6a+9 }=\frac{ 4(a^2-a-6) }{ a^2-6a+9 }=\frac{ 4(a-3)(a+2) }{ (a-3)(a-3) }=\frac{ 4(a+2) }{ (a-3) }\]\[\frac{ 8 }{ (9-a^2) }=\frac{ 8 }{ (3+a)(3-a) }\]So now we have\[\frac{ 8 }{ (3+a)(3-a) }\div\frac{ 4(a+2) }{ (a-3) }=\frac{ 8 }{ (3+a)(3-a) }\times\frac{ (a-3) }{ 4(a+2) }=\frac{ 8 (a-3)}{ (3+a)(3-a)4(a+2)}\]I think distributing will be the last step. I'll let you do that.\[\frac{ 2 (a-3)}{ (3+a)(3-a)(a+2)}\]

OpenStudy (anonymous):

(a+2) was cut off